OpenStudy (anonymous):

What is 3g^2 - 16gh -12h^2 in factored form? PLEASE HELP ME

OpenStudy (mathstudent55):

First, just look at the variables. g^2 = g*g, and h^2 = h*h From the first and last steps of FOIL, where would g and h go in the two binomials that this will factor into?

OpenStudy (anonymous):

i have no idea -__-

OpenStudy (mathstudent55):

When you use FOIL, this is how it works: (a + b)(c + d) = ac + ad + bc + bd Look at the first term in the result, ac. It comes from multiplying the two first terms together. In your problem you have 3g^2, as your first term. 3g^2 comes from multiplying together 3g and g. That means the first step to factoring this is (3g )(g )

OpenStudy (mathstudent55):

The last term of your question is -12h^2. That comes from multiplying two terms with h in them. You'll get something like this: (3g h)(g h) Now you need to figure out what factors with h will multiply to -12h^2

OpenStudy (anonymous):

yea ik that but where does the -16gh come into play??

OpenStudy (mathstudent55):

That comes from multiplying the two terms OI in FOIL.

OpenStudy (mathstudent55):

To get -12h^, there are several possibilities: -12 = -12 * 1 -12 = -1 * 12 -12 = -3 * 4 -12 = -4 * 3 -12 = -2 * 6 -12 = -6 * 2 There are many combinations of numbers that will give you a product of -12 as shown above, but you need to choose the correct one for this problem.

OpenStudy (mathstudent55):

Let's say you choose -4 and 3. Then you'd have (3g - 4h)(g + 3h) Now foil it and see if you get your original problem. (3g - 4h)(g + 3h) = 3g^2 + 9gh - 4gh - 12h^2 = 3g^2 + 5gh - 12h^2 Since we get a middle term of -5gh and the problem has a middle term of -16gh, our guess didn't work.

OpenStudy (mathstudent55):

Let's try -6 and 2 (3g + 2h)(g - 6h) = 3g^2 - 18gh + 2gh - 12h^2 = 3g^2 - 16gh - 12h^2 Which is the problem you started with so this is the answer:. (3g + 2h)(g - 6h)

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct